# Photocathode characterisation for robust PICOSEC Micromegas precise-timing detectors

### v1

@@ 4.1 Cesium Iodide
Cesium Iodide is the most widely used material for photocathodes in gaseous radiation detectors due to its high QE and its sensitivity to UV radiation. The baseline photocathode used in the PICOSEC detector consists of an 18 nm thick CsI layer deposited onto a 3 mm thick MgF2 substrate with a 3 nm thick Cr conductive interfacial layer, exceeding 12 photoelectrons per MIP [1], Previous studies on transparency in the VUV range have shown that the pure MgF2 substrate has approximately 80% transparency, which decreases to 45% with the addition of the Cr layer, and further to 10% with the CsI layer [5].
Particle beam measurements were performed using a PICOSEC prototype equipped with a CsI photocathode, operated in a sealed mode at a gas pressure of 985 mbar. At voltage settings of V = -430 V on the cathode and V = 265 V on the anode, the detector exhibited a time resolution of = 15.8 2.2 ps, as illustrated in Fig. 5.
Despite exhibiting excellent time resolution, CsI is vulnerable to damage from ion backflow, discharges and humidity. Fig. 6 illustrates the decline in QE observed in a CsI sample before and after multiple irradiation steps, providing evidence of ion bombardment affecting CsI. Studies demonstrated that after accumulating a charge of 6 mC/cm2, the QE for CsI decreased by 80%, whereas for B4C, it decreased by 45%, indicating a more rapid degradation of CsI [5]. One possible solution to reduce the impact of ion backflow on CsI and mitigate degradation involves introducing protective layers, such as MgF2 or Lithium Fluoride (LiF) [18]. Nonetheless, tests conducted within the PICOSEC project revealed that these layers inhibited electron extraction, leading to reduced QE [5]. Another approach to address the issue of non-robustness is to explore alternative materials, with carbon-based photocathodes like DLC and B4C being the most promising candidates.
@@ 4.2 Diamond-Like Carbon
Diamond-Like Carbon represents an alternative photocathode material that offers greater resistance to environmental influences compared to CsI. DLC belongs to a class of amorphous carbon materials, exhibiting excellent properties including chemical and thermal stability, hardness and robustness. Initial measurements of DLC photocathodes have been conducted by collaborators from USTC [13], including tests of various layer thicknesses, measurements of QE, NPE and time resolution as well as aging studies. DLC samples consistently demonstrated good performance and robustness, making them a strong candidate for PICOSEC photocathodes.
@@
The DLC photocathodes presented in this paper were fabricated using a magnetron sputtering technique at the CERN Micro-Pattern Technologies (MPT) workshop. DLC photocathodes with thicknesses ranging from 1 nm to 100 nm were deposited on glass and MgF2 substrates, with some samples including a Cr interfacial layer. Fig. 7 depicts the DLC photocathodes with the various thicknesses produced during one of the two deposition campaigns. The process proved to be reproducible across both campaigns.
@@
Profilometer measurements were performed to investigate the thicknesses of the 50 nm and 100 nm layers. For layers in the range of a few nanometers, specifically from 1.5 nm to 4.5 nm, thicknesses were estimated by scaling the coating time. Fig. 8 presents the results of transparency measurements conducted in the wavelength range from 200 nm to 800 nm using a spectrophotometer. The data demonstrate a correlation between the estimated thicknesses and the measured transparency. The transparency of samples with a thickness of a few nanometers in the VUV range is approximately 60%.
Surface resistivity measurements were conducted, exhibiting a correlation between the thickness of the DLC layer and its resistivity, as shown in Fig. 9. The higher resistivity of the DLC photocathodes on the MgF2 radiator compared to the glass substrate suggests that the layer deposited on the MgF2 is thinner, potentially due to the lower adhesion of the crystal.
@@
The DLC photocathodes, with thicknesses ranging from 1.5 nm to 3.5 nm, were characterised during particle beam measurements. The measurements were performed at a gas pressure of 990 mbar. The NPE generated by the DLC photocathodes varied between 2.5 and 3. The 10 mm diameter active area detector featuring a 1.5 nm thick DLC photocathode deposited directly on the radiator and operated at voltages of V = -500 V on the cathode and V = 275 V on the anode demonstrated the best performance, achieving a time resolution of = 31.9 1.3 ps, as illustrated in Fig. 10. Thicker samples exhibited comparable time resolution values, with a deviation of approximately 4 ps.
@@
To enhance UV photon production, a thicker radiator was introduced. At the same time, the Cherenkov cone diameter was widened. A prototype with a 15 mm diameter active area, equipped with a 2.5 nm thick DLC photocathode deposited on a 5 mm thick MgF2 radiator, was tested. The analysis involved applying a geometric cut, consisting of a 5 mm diameter circle around the pad center, to include only fully contained events. The detector operated at of V = -490 V on the cathode and V = 275 V on the anode yielded a time resolution of = 28.0 1.4 ps.
@@ 4.3 Boron Carbide
The B4C photocathodes were evaluated through particle beam measurements. The prototypes were operated at a gas pressure of 990 mbar. The NPE created by the B4C photocathodes ranged between 2 and 4, with higher PE production observed in thinner layers. The detector equipped with a 9 nm thick B4C photocathode and operated at V = -490 V on the cathode and V = 275 V on the anode, exhibited the best performance, achieving a time resolution of = 34.5 1.5 ps, as shown in Fig. 11. The time resolution of the remaining B4C samples decrease, with the thickest sample showing a degradation of approximately 10 ps.
@@ 5 Discussion
In the comparative analysis of photocathode materials, distinct performance characteristics were observed among CsI, DLC, and B4C. CsI exhibited the highest QE, exceeding 12 photoelectrons per MIP, whereas DLC yielded QE values ranging between 2.5 and 3, while B4C between 2 and 4. Despite B4C being more transparent than DLC, both materials exhibit similar efficiency, generating comparable NPE. NPE correlates with the achievable time resolution. CsI showcased the best performance with 15.8 ps, followed by DLC with 32 ps and B4C with 34.5 ps. Although demonstrating excellent timing results, CsI also exhibited a faster degradation in QE compared to carbon-based materials. Furthermore, carbon-based materials displayed notably greater resistance to humidity. Considering the demand for robust detectors with timing specifications that allow a margin on operating condition, DLC and B4C emerge as promising candidates. Consequently, a detector configuration comprising a single-pad prototype with a 15 mm diameter active area resistive Micromegas of 20 M/, utilising a 1.5 nm DLC photocathode and an amplifier integrated on the outer PCB, was tested. Operating the device at V = -530 V on the cathode and V = 275 V on the anode, a time resolution of = 31.4 0.6 ps was achieved within a 9 mm diameter circle around the pad center, including only fully contained events. A uniform time response across this region, with an RMS = 38.8 0.3 ps, was observed, as depicted in Fig. 12.
@@ 6 Conclusions
The work described in this paper is focused on enhancing the robustness of PICOSEC detectors. The research delves into the characterisation of robust carbon-based photocathodes, such as DLC and B4C. Results obtained from prototypes equipped with DLC and B4C photocathodes revealed time resolutions of approximately 32 ps and 34.5 ps, correspondingly. Efforts dedicated to detector developments increase the feasibility of the PICOSEC concept for experiments requiring sustained performance while maintaining excellent timing precision. Current developments include research on alternative materials, such as titanium to replace Cr, and carbon-based nanostructures for use as photocathodes. Simultaneously, a 10×10 cm2 robust photocathode, incorporating a conductive interlayer to prevent a voltage drop, will be tested with a 100-channel prototype [10] and a SAMPIC digitiser [19]. In view of improving the stability of the detector, the production of a high-rate 10×10 cm2 Micromegas with double-layer DLC for vertical charge evacuation and evaluation of rate capability is ongoing. Additionally, efforts to enhance spatial resolution involve testing high-granularity layouts. Scaling up the PICOSEC detector by tiling 10×10 cm2 modules or developing larger prototypes are the next steps.

### v2

@@ 4.1 Cesium Iodide
Cesium Iodide is the most widely used material for photocathodes in gaseous radiation detectors due to its high QE and sensitivity to UV radiation. Notable applications of CsI as a UV-to-electron converter include its use in the RICH detectors of the particle identification systems for ALICE [12, 13] and COMPASS [22]. The baseline photocathode used in the PICOSEC detector consists of an 18 nm thick CsI layer deposited onto a 3 mm thick MgF2 substrate with a 3 nm thick Cr conductive interfacial layer, exceeding 12 photoelectrons per MIP [2]. Previous studies on transparency in the VUV range have shown that the pure MgF2 substrate has approximately 80% transparency, which decreases to 45% with the addition of the Cr layer, and further to 10% with the CsI layer [5].
Particle beam measurements were performed using a PICOSEC prototype equipped with a CsI photocathode, operated in a sealed mode at a gas pressure of 985 mbar. At voltage settings of V = -430 V on the cathode and V = 265 V on the anode, the detector exhibited a time resolution of = 15.8 2.2 ps, as illustrated in Fig. 6, with a full detection efficiency for completely contained events.
Despite exhibiting excellent time resolution, CsI is vulnerable to damage from ion backflow and discharges. Figure 7 illustrates the decline in QE observed in a CsI sample before and after multiple irradiation steps, providing evidence of ion bombardment affecting CsI. Studies demonstrated that after accumulating a charge of 6 mC/cm2, the QE for CsI decreased by 80%, whereas for B4C, it decreased by 45%, indicating a more rapid degradation of CsI [5]. One possible solution to reduce the impact of ion backflow on CsI and mitigate degradation involves introducing protective layers, such as MgF2 or Lithium Fluoride (LiF) [23]. Nonetheless, tests conducted within the PICOSEC project revealed that these layers inhibited electron extraction, leading to reduced QE [5]. Additionally, CsI photocathodes are sensitive to humidity and require storage in a vacuum or dry gas to maintain good time resolution, as they degrade within minutes of humidity exposure. To address the issue of non-robustness, one approach is to explore alternative materials, with carbon-based photocathodes such as DLC and B4C emerging as the most promising candidates.
@@ 4.2 Diamond-Like Carbon
Diamond-Like Carbon represents an alternative photocathode material that offers greater resistance to environmental influences compared to CsI. DLC belongs to a class of amorphous carbon materials, exhibiting excellent properties including chemical and thermal stability, hardness and robustness. Initial measurements of DLC photocathodes have been conducted by collaborators from USTC [7], including tests of various layer thicknesses, measurements of QE, and time resolution as well as aging studies. DLC samples consistently demonstrated good performance and robustness, making them a strong candidate for PICOSEC photocathodes.
@@
The DLC photocathodes presented in this paper were fabricated using a magnetron sputtering technique at the CERN Micro-Pattern Technologies (MPT) workshop. DLC photocathodes with thicknesses ranging from 1 nm to 100 nm were deposited on glass and MgF2 substrates, with some samples including a Cr interfacial layer. The process was reproducible across both campaigns.
@@
Profilometer measurements were performed to investigate the thicknesses of the 50 nm and 100 nm layers. For layers in the range of a few nanometers, specifically from 1.5 nm to 4.5 nm, thicknesses were estimated by scaling the coating time. Figure 8 presents the results of transparency measurements conducted in the wavelength range from 200 nm to 800 nm using a spectrophotometer. The data demonstrate a correlation between the estimated thicknesses and the measured transparency. The DLC photocathodes deposited on both substrates exhibit similar behaviour; nevertheless, the glass samples show slightly reduced transparency and a low-wavelength cutoff below 300 nm. The transparency of MgF2 samples with a thickness of a few nanometers at 160 nm is about 60%.
Surface resistivity measurements were conducted as an additional test to confirm whether the estimated thicknesses follow the expected trend, where thinner samples exhibit higher surface resistivity. The measurements were performed using the picoammeter [16]. A low voltage was applied across two conductive strips coated on the substrates and the resistivity of the photocathodes was recorded. The results exhibit a correlation between the thickness of the DLC layer and its resistivity, as shown in Fig. 9. The surface resistivity of the 1.5 nm DLC samples deviates from the expected trend, which may be attributed to the potential non-uniformity of these very thin layers. Additionally, the higher values for the DLC photocathodes on the MgF2 radiator compared to the glass substrate suggests that the layer deposited on the MgF2 is thinner, potentially due to the lower adhesion of the crystal.
@@
The DLC photocathodes with thicknesses ranging from 1.5 nm to 3.5 nm were characterised during particle beam measurements. The measurements were performed at a gas pressure of 990 mbar. The generated by the DLC photocathodes varied between 2.5 and 3. The 10 mm diameter active area detector featuring a 1.5 nm thick DLC photocathode deposited directly on the radiator and operated at voltages of V = -500 V on the cathode and V = 275 V on the anode demonstrated the best performance, achieving a time resolution of = 31.9 1.3 ps, as illustrated in Fig. 10, with a detection efficiency of 96.8%. Thicker samples exhibited comparable time resolution values, with approximately 4 ps worse performance.
@@
To enhance UV photon production, a thicker radiator was introduced. At the same time, the Cherenkov cone diameter was widened. A prototype with a 15 mm diameter active area, equipped with a 2.5 nm thick DLC photocathode deposited on a 5 mm thick MgF2 radiator, was tested. The analysis involved applying a geometric cut, consisting of a 5 mm diameter circle around the pad center, to include only fully contained events. The detector operated at of V = -490 V on the cathode and V = 275 V on the anode yielded a time resolution of = 28.0 1.4 ps, with a 99.4% detection efficiency.
@@ 4.3 Boron Carbide
The B4C photocathodes were evaluated through particle beam measurements. The prototypes were operated at a gas pressure of 990 mbar. The created by the B4C photocathodes ranged between 2 and 4, with higher photoelectron production observed for thinner layers. The detector equipped with a 9 nm thick B4C photocathode and operated at V = -490 V on the cathode and V = 275 V on the anode, exhibited the best performance, achieving a time resolution of = 34.5 1.5 ps, as shown in Fig. 11, with a 95.4% detection efficiency. Thicker B4C samples showed worse time resolutions, with the thickest sample by around 10 ps.
@@ 5 Discussion
In the comparative analysis of photocathode materials, distinct performance characteristics were observed among CsI, DLC, and B4C. CsI exhibited the highest , exceeding 12 photoelectrons per MIP, whereas DLC yielded ranging between 2.5 and 3, while B4C between 2 and 4. Despite B4C being more transparent than DLC, both materials exhibit similar efficiency, generating comparable , which correlates with the achievable time resolutions. CsI showed the best performance with 15.8 ps, followed by DLC with 32 ps and B4C with 34.5 ps. Although CsI achieved excellent time resolution, it showed faster degradation in QE due to ion backflow, compared to carbon-based materials. Additionally, the carbon-based photocathodes are not sensitive to humidity and have been successfully stored in air without any visual or performance degradation. These samples have been reused across multiple test beam campaigns over several years, demonstrating no deterioration in time resolution.
Considering the demand for robust detectors with timing specifications that allow a margin on operating condition, DLC and B4C emerge as promising candidates. Consequently, a detector configuration comprising a single-pad prototype with a 15 mm diameter active area resistive Micromegas of 20 M/, utilising a 1.5 nm DLC photocathode and an amplifier integrated on the outer PCB, was tested. Operating the device at V = -530 V on the cathode and V = 275 V on the anode, a time resolution of = 31.4 0.6 ps was achieved within a 9 mm diameter circle around the pad center, including only fully contained events. A good time response across this region, with an RMS = 38.8 0.3 ps, was observed, as depicted in Fig. 12. A detection efficiency was around 97.6%.
@@ 6 Conclusions
The work described in this paper focuses on enhancing the robustness of PICOSEC detectors. The research delves into the characterisation of carbon-based photocathodes, such as DLC and B4C. Results obtained from prototypes equipped with these photocathodes showed time resolutions of approximately 32 ps and 34.5 ps, respectively. Furthermore, carbon-based materials displayed notably greater resistance to ion backflow, discharges and humidity. The efforts dedicated to advancing the detectors increase the feasibility of the PICOSEC concept for experiments requiring sustained performance while maintaining excellent timing precision.
Current PICOSEC developments include research on alternative materials, such as titanium to replace Cr, and carbon-based nanostructures for use as photocathodes. Regarding the robustness of photocathodes under different humidity conditions, specific measurements are planned. Simultaneously, a 10×10 cm2 robust photocathode, incorporating a conductive interlayer to prevent a voltage drop, will be tested with a 100-channel prototype [9, 10] and a SAMPIC digitiser [24]. In view of improving the stability of the detector for high-rate conditions, the production of a 10×10 cm2 Micromegas with double-layer DLC for vertical charge evacuation and evaluation of rate capability is ongoing. Additionally, efforts to enhance spatial resolution involve testing high-granularity layouts. Scaling up the PICOSEC detector by tiling 10×10 cm2 modules or developing larger prototypes are the next steps.